# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v1

@@ 3 Metrics of Hyperbolic LVMs
Tosi et al., (2014) introduced the pullback metric for Euclidean GPLVMs and its use to compute geodesics that adhere to the data distribution. Here, we extend this approach to hyperbolic latent variable models, specifically the GPHLVM introduced in Sec. 2. We show that the resulting metric can be computed similarly to the Euclidean case, with two key exceptions: (1) the derivatives of the GPHLVM nonlinear mapping must be projected onto appropriate tangent spaces; and (2) the computation of the derivatives for the hyperbolic heat kernels is significantly more challenging than in the Euclidean scenario.
3.1 Hyperbolic Pullback Metric
For a deterministic mapping $f:\mathbb{R}^{D_{x}}\rightarrow\mathbb{R}^{D_{y}}$, the pullback metric is defined as $\bm{G}_{\bm{x}^{*}}=\bm{J}^{\mathsf{T}}\bm{J}$, where $\bm{J}\in\mathbb{R}^{D_{y}\times D_{x}}$ is the Jacobian of $f$ at $\bm{x}^{*}$. In the context of LVMs, the mapping $f$ is stochastic. As Tosi et al., (2014), we consider LVMs for which the probability over $\bm{J}$ follows a Gaussian distribution. Specifically, assuming independent rows $\bm{J}_{d}\in\mathbb{R}^{D_{x}}$, each with its own mean but shared covariance matrix, the Jacobian distribution is of the form,
@@ 3.1 Hyperbolic Pullback Metric
Therefore, the metric tensor $\bm{G}_{\bm{x}^{*}}$ follows a non-central Wishart distribution (Anderson,, 1946),
$p(\bm{G}_{\bm{x}^{*}})=\mathcal{\bm{W}}_{D_{x}}(D_{y},\bm{\Sigma}_{J},\mathbb{E}[\bm{J}]^{\mathsf{T}}\mathbb{E}[\bm{J}])\,,$ (8)
@@
$\mathbb{E}[\bm{G}_{\bm{x}^{*}}]=\mathbb{E}[\bm{J}]^{\mathsf{T}}\mathbb{E}[\bm{J}]+D_{y}\bm{\Sigma}_{J}\,.$ (9)
While the hyperbolic case follows a similar strategy, special care is required to ensure that the Jacobian rows lie on appropriate tangent spaces. Starting from the mapping $f:\mathbb{H}^{D_{x}}_{\mathcal{L}}\rightarrow\mathbb{R}^{D_{y}}$, the Riemannian Jacobian is composed by the Riemannian gradients for each output dimension, i.e., ${\tilde{\bm{J}}=[\mathrm{grad}_{\bm{x}}(f_{1})\>\ldots\>\mathrm{grad}_{\bm{x}}(f_{D_{y}})]^{\mathsf{T}}}$. As for Riemannian submanifolds (Boumal,, 2023), the Lorentz gradient ${\mathrm{grad}_{\bm{x}}(f)\in\mathcal{T}_{\bm{x}}\mathbb{H}^{D_{x}}_{\mathcal{L}}}$ equals the orthogonal projectiong of the Euclidean gradient $\frac{\partial f}{\partial\bm{x}}$ onto the tangent space,
$\mathrm{grad}_{\bm{x}}(f)=\mathrm{proj}_{\bm{x}}\left(\frac{\partial f}{\partial\bm{x}}\right)=\bm{P}_{\bm{x}}\frac{\partial f}{\partial\bm{x}},$ (10)
with $\mathrm{proj}_{\bm{x}}:\mathbb{R}^{D_{x}+1}\to\mathcal{T}_{\bm{x}}\mathbb{H}^{D_{x}}_{\mathcal{L}}$ (see App. A for details). Therefore, the Riemannian Jacobian $\tilde{\bm{J}}$ is computed from the Euclidean Jacobian $\bm{J}$ as $\tilde{\bm{J}}=\bm{J}\bm{P}^{\mathsf{T}}_{\bm{x}^{*}}$, and follows the Jacobian distribution,
$p(\tilde{\bm{J}})=\prod_{d=1}^{D_{y}}\mathcal{N}(\tilde{\bm{J}}_{d}\mid\bm{P}_{\bm{x}^{*}}\bm{\mu}_{d},\bm{P}_{\bm{x}^{*}}\bm{\Sigma}_{J}\bm{P}_{\bm{x}^{*}}^{\mathsf{T}})\,.$ (11)
Similarly to (9), the expected pullback metric tensor follows a non-central Wishart distribution,
$\mathbb{E}[\bm{G}_{\bm{x}^{*}}]$ $=\mathbb{E}[\tilde{\bm{J}}]^{\mathsf{T}}\mathbb{E}[\tilde{\bm{J}}]+D_{y}\tilde{\bm{\Sigma}}_{J}$
$=\bm{P}_{\bm{x}^{*}}(\mathbb{E}[\bm{J}]^{\mathsf{T}}\mathbb{E}[\bm{J}]+D_{y}\bm{\Sigma}_{J})\bm{P}_{\bm{x}^{*}}^{\mathsf{T}}\,.$ (12)
3.2 GPHLVM Pullback Metric
Here, we consider the case where the mapping $f$ is defined as a GPHLVM. We derive the distribution of the Riemannian Jacobian, which we then use to compute the expected GPHLVM pullback metric $\bm{G}_{\bm{x}^{*}}$. As for GPLVMs (Tosi et al.,, 2014), the joint distribution of the Jacobian $\tilde{\bm{J}}$ and observations $\bm{Y}$ is given as,
@@ 3.2 GPHLVM Pullback Metric
where we rely on the hyperbolic kernel derivatives
$\partial k(\bm{X},\bm{x}^{*})$ $=\frac{\partial}{\partial\bm{x}^{\prime}}k(\bm{X},\bm{x}^{\prime})\Big{|}_{\bm{x}^{\prime}=\bm{x}^{*}},$ (14)
$\partial^{2}k(\bm{x}^{*},\bm{x}^{*})$ $=\frac{\partial^{2}}{\partial\bm{x}^{\prime}\,\partial\bm{x}}k(\bm{x},\bm{x}^{\prime})\Big{|}_{\bm{x}=\bm{x}^{\prime}=\bm{x}^{*}}\,.$ (15)
While in the Euclidean case the kernel derivatives are straightforward, the hyperbolic setting is more challenging, as will be discussed in Sec. 4. Conditioning on the observations $\bm{Y}_{d}\in\mathbb{R}^{N}$ results in the probability distribution of the Jacobian at $\bm{x}^{*}$ (Bishop,, 2006),
@@
Using Equation (12) we compute the final expected pullback metric tensor on the tangent space of $\bm{x}^{*}$ as,
$\mathbb{E}[\bm{G}_{\bm{x}^{*}}]=\bm{P}_{\bm{x}^{*}}(\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J})\bm{P}_{\bm{x}^{*}}^{\mathsf{T}}\,,$ (20)
@@
3.3 Pullback Geodesics
Having endowed the hyperbolic latent space with a pullback metric, we are now interested in computing shortest paths, i.e., geodesics, between latent hyperbolic points. While hyperbolic geodesics have a closed-form solution, they may not always align with the underlying data distribution (see Fig. 1). Instead, pullback geodesics naturally follow the data distribution due to the pullback metric, but require solving an optimization problem. Specifically, pullback geodesics are computed by minimizing the curve length, or equivalently the curve energy $E$ with respect to the pullback metric. When considering a discretized geodesic composed by a set of $M$ points $\bm{x}_{i}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$, this boils down to minimize,
$E=\sum_{i=0}^{M-2}\bm{v}_{i}^{\mathsf{T}}\bm{G}_{\bm{x}_{i}}\bm{v}_{i}\,,\quad\text{with}\quad\bm{v}_{i}=\text{Log}_{\bm{x}_{i}}({\bm{x}_{i+1}}).$ (21)
While it is possible to iteratively optimize the curve points $\bm{x}_{i}$ directly, this often leads to uneven spacing among them. To address this issue, one could either use a parametric curve on the manifold (Gousenbourger et al.,, 2014) and optimize its parameters instead of the points directly; or alternatively, incorporate the spline energy as a regularization factor in the optimization process (Heeren et al.,, 2018). The latter is the approach we follow. Specifically, the spline energy is given as $E_{\text{spline}}\approx\sum_{i=1}^{M-2}\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x}_{i},\bar{\bm{x}}_{i})^{2}$, where $\bar{\bm{x}}_{i}=\text{Exp}_{\bm{x}_{i-1}}\left(\frac{1}{2}\,\text{Log}_{\bm{x}_{i-1}}(\bm{x}_{i+1})\right)$ is the geodesic midpoint between $\bm{x}_{i-1}$ and $\bm{x}_{i+1}$. The final optimization problem is,
@@ 3.3 Pullback Geodesics
with $\lambda$ weighting the influence of the regularization. Note that the optimization parameters are Riemannian, so we leverage Riemannian optimizers such as Riemannian Adam (Bécigneul and Ganea,, 2019) to optimize (21).
@@ 4 Hyperbolic Kernel Derivatives
In this section, we discuss the challenges encountered when using autodiff tools, in particular PyTorch  (Paszke et al.,, 2019), to compute the derivatives of the hyperbolic SE kernels. We observed that, as reported in Table 1, autodiff-based kernel derivatives are approximately five times slower than custom implementations in the two-dimensional case. More importantly, in the three-dimensional case, autodiff fails to compute the derivatives entirely.
2D kernel ($L=3000)$ 3D kernel
@@
Table 1: Average computation times in seconds over $100$ runs, each evaluating the pullback metric tensor and its derivatives at a random point. In the 3D case computing kernel derivatives with PyTorch is not feasible.
4.1 2D Hyperbolic Heat Kernel Derivatives
In order to compute the 2D hyperbolic heat kernel $k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{z})$ [^1] We use $\bm{z}$ to denote the second input and avoid using a prime symbol. we rely on the Monte Carlo approximation given in Equation (4). To compute the pullback metric tensor $\bm{G}_{\bm{x}^{*}}$ at a point $\bm{x}^{*}$ we need the kernel derivatives $\frac{\partial}{\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{x}^{*}}$ and $\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$, see Equations (14) and (15). The only part of the kernel (6) that depends on input points is the helper function $\Phi_{l}(\bm{x},\bm{z})$, whose derivatives are given by,
@@
Using these derivatives we can compute the pullback metric tensor itself. Minimizing the curve energy (21), however, requires the derivative of the pullback metric tensor $\frac{\partial}{\partial\bm{x}^{*}}\bm{G}_{\bm{x}^{*}}^{\mathcal{L}}$. This, in turn, requires the derivative of the Jacobian of the covariance matrix $\frac{\partial}{\partial\bm{x}^{*}}\bm{\Sigma}_{J}$, which depends on the kernel derivatives $\frac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$, and $\frac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$. Again the derivatives are effectively determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$,
$\!\frac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}\Phi_{l}(\bm{x},\bm{z})\!$ $=\!\left[\!\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\phi_{l}(\bm{x}_{\mathcal{P}}\!)\!\right]\!\times_{1}\!\left[\!\dfrac{\partial}{\partial\bm{z}}\bar{\phi}_{l}(\bm{z}_{\mathcal{P}}\!)\!\right]^{\mathsf{T}}\!\!\!,$ (25)
$\!\frac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}\Phi_{l}(\bm{x},\bm{z})\!$ $=\!\left[\!\dfrac{\partial^{2}}{\partial\bm{z}^{2}}\bar{\phi}_{l}(\bm{z}_{\mathcal{P}}\!)\!\right]\!\times_{0}\!\left[\!\dfrac{\partial}{\partial\bm{x}}\phi_{l}(\bm{x}_{\mathcal{P}}\!)\!\right]^{\mathsf{T}}\!\!\!,$ (26)
@@
where $\bm{A}\times_{i}\bm{B}$ refers to the tensor product with the $i$-th dimension of tensor $\bm{A}$ being multiplied with the first dimension of matrix $\bm{B}$. Additionally, if matrix $\bm{B}$ is a row vector, ×$\times_{i}$ indicates that the tensor is unsqueezed at dimension $i$ before performing the multiplication. In Equations (25) and (26) the matrices $\frac{\partial^{2}}{\partial\bm{x}^{2}}\phi_{l}(\bm{x}_{\mathcal{P}}),\frac{\partial^{2}}{\partial\bm{z}^{2}}\phi_{l}(\bm{z}_{\mathcal{P}})\in\mathbb{R}^{3\times 3}$ are therefore effectively interpreted as $\mathbb{R}^{3\times 1\times 3}$ and $\mathbb{R}^{1\times 3\times 3}$, respectively. The complete derivation of all kernel derivatives is provided in App. B.
Figure 1: Left and middle: Euclidean and hyperbolic pullback metrics on a $\mathsf{C}$-shape trajectory ( ). The geodesic of the base manifold ( ) and the geodesic optimized on the pullback metric ( ) are represented on both Euclidean and hyperbolic spaces. Right: Curve energy along the geodesics on the Euclidean (top) and hyperbolic (bottom) cases. These include the base manifold geodesic ( ), the base manifold geodesic with energy (21) evaluated using the pullback metric ( ), and the geodesic optimized with respect to the pullback metric ( ).
4.2 3D Hyperbolic Heat Kernel Derivatives
@@
In the three-dimensional case, the hyperbolic heat kernel is given as $k^{\mathbb{H}^{3}_{\mathcal{L}}}(\bm{x},\bm{z})=\frac{\rho}{s}e^{-\frac{\rho^{2}}{\kappa}}$. Note that, compared to Equation (5), we simplified the notation by omitting the variance $\tau$ and normalization constant $C_{\infty}$, by merging $2\kappa^{2}$ into just $\kappa$, and by introducing helper variables $u=\langle\bm{x},\bm{z}\rangle_{\mathcal{L}}$, and $s=\sqrt{u^{2}-1}$. As in the 2D case, we compute the first two kernel derivatives $\frac{\partial}{\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{x}^{*}}$ and $\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$ in order to compute the pullback metric tensor $\bm{G}_{\bm{x}^{*}}$, as follows,
$\dfrac{\partial}{\partial\bm{x}}\,k(\bm{x},\bm{z})$ $=g(u)\,\bm{G}\bm{z}\,e^{-\frac{\rho^{2}}{\kappa}}\,,$ (27)
$\dfrac{\partial^{2}}{\partial\bm{z}\partial{\bm{x}}}k(\bm{x},\bm{z})$ $=\left(h(u)\bm{G}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}+g(u)\bm{G}\right)e^{-\frac{\rho^{2}}{\kappa}}\,,$ (28)

### v2

@@ 3 Metrics of Hyperbolic LVMs
Tosi et al. (2014) introduced the pullback metric for Euclidean GPLVMs and its use to compute geodesics that adhere to the data distribution. Here, we extend this approach to hyperbolic LVMs, specifically to the GPHLVM introduced in Sec. 2. We show that the resulting metric can be computed similarly to the Euclidean case, with two key exceptions: (1) the pullback metric must be adequately projected onto appropriate tangent spaces; and (2) the computation of the derivatives for the hyperbolic SE kernels is significantly more challenging than in the Euclidean scenario.
3.1 A General Hyperbolic Pullback Metric
A deterministic immersion ${f:\mathbb{R}^{D_{x}}\to\mathbb{R}^{D_{y}}}$ pulls back the metric into the latent space following (7), i.e., ${\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}=\bm{J}^{\mathsf{T}}\bm{J}}$. In the context of LVMs, the immersion $f$ is stochastic. As developed in (Tosi et al., 2014), the stochastic immersion $f$ induces a distribution over its Jacobian $\bm{J}$, which itself induces a distribution over the pullback metric. As Tosi et al. (2014), we consider LVMs for which the probability over $\bm{J}$ follows a Gaussian distribution. Assuming independent rows $\bm{J}_{d}\in\mathbb{R}^{D_{x}}$, each with its own mean but shared covariance matrix, the Jacobian distribution is of the form,
@@ 3.1 A General Hyperbolic Pullback Metric
Therefore, the metric tensor $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ follows a non-central Wishart distribution (Anderson, 1946),
$p(\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}})=\mathcal{\bm{W}}_{D_{x}}(D_{y},\bm{\Sigma}_{J},\mathbb{E}[\bm{J}]^{\mathsf{T}}\mathbb{E}[\bm{J}])\,,$ (9)
@@
$\mathbb{E}[\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}]=\mathbb{E}[\bm{J}]^{\mathsf{T}}\mathbb{E}[\bm{J}]+D_{y}\bm{\Sigma}_{J}\,.$ (10)
While the hyperbolic case follows a similar strategy, special care is required to ensure that the Jacobian rows lie on appropriate tangent spaces. A hyperbolic LVM defines a stochastic mapping $f:\mathbb{H}^{D_{x}}_{\mathcal{L}}\rightarrow\mathbb{R}^{D_{y}}$, whose Jacobian is computed from the Riemannian gradients of $f$ as in (1). As for Riemannian submanifolds (Boumal, 2023), each Riemannian gradient $\mathrm{grad}_{\bm{x}}(f_{d})$ equals the orthogonal projection of the Euclidean gradient $\frac{\partial f_{d}}{\partial\bm{x}}\in\mathbb{R}^{D_{x}+1}$ onto $\mathcal{T}_{\bm{x}}\mathbb{H}^{D_{x}}_{\mathcal{L}}$,
$\mathrm{grad}_{\bm{x}}(f_{d})=\mathrm{proj}_{\bm{x}}\left(\frac{\partial f_{d}}{\partial\bm{x}}\right)=\bm{P}_{\bm{x}}\frac{\partial f_{d}}{\partial\bm{x}},$ (11)
with $\bm{P}_{\bm{x}}\!=\!\bm{G}^{\mathcal{L}}+\bm{x}\bm{x}^{\mathsf{T}}$. Therefore, the Riemannian Jacobian $\tilde{\bm{J}}$ is computed from the Euclidean Jacobian $\bm{J}$ as ${\tilde{\bm{J}}\!=\![\bm{P}_{\bm{x}}\frac{\partial f_{1}}{\partial\bm{x}}\>\ldots\>\bm{P}_{\bm{x}}\frac{\partial f_{D_{y}}}{\partial\bm{x}}]^{\mathsf{T}}\!=\!\bm{J}\bm{P}^{\mathsf{T}}_{\bm{x}}}$. Assuming independent rows $\tilde{\bm{J}}_{d}$ as in the Euclidean case, the Jacobian distribution is computed from (8) as (Bishop, 2006, Chap. 2.3),
$p(\tilde{\bm{J}})=\prod_{d=1}^{D_{y}}\mathcal{N}(\tilde{\bm{J}}_{d}\mid\bm{P}_{\bm{x}}\bm{\mu}_{d},\bm{P}_{\bm{x}}\bm{\Sigma}_{J}\bm{P}_{\bm{x}}^{\mathsf{T}})\,.$ (12)
Similarly to $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ in (9), the metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ follows a non-central Wishart distribution (Anderson, 1946),
$p(\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}})=\mathcal{\bm{W}}_{D_{x}}(D_{y},\tilde{\bm{\Sigma}}_{J},\mathbb{E}[\tilde{\bm{J}}]^{\mathsf{T}}\mathbb{E}[\tilde{\bm{J}}])\,,$ (13)
with $\tilde{\bm{\Sigma}}_{J}\!=\!\bm{P}_{\bm{x}}\bm{\Sigma}_{J}\bm{P}_{\bm{x}}^{\mathsf{T}}$, leading to the expected metric,
$\mathbb{E}[\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}]$ $=\mathbb{E}[\tilde{\bm{J}}]^{\mathsf{T}}\mathbb{E}[\tilde{\bm{J}}]+D_{y}\tilde{\bm{\Sigma}}_{J}$
$=\bm{P}_{\bm{x}}(\mathbb{E}[\bm{J}]^{\mathsf{T}}\mathbb{E}[\bm{J}]+D_{y}\bm{\Sigma}_{J})\bm{P}_{\bm{x}}^{\mathsf{T}}\,.$ (14)
Intuitively, the pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ is the orthogonal projection of the Euclidean pullback metric $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ onto the tangent space $\mathcal{T}_{\bm{x}}\mathbb{H}^{D_{x}}_{\mathcal{L}}$ (see Fig. 1). It relates to the hyperbolic metric $\bm{G}^{\mathcal{L}}$ via the projection $\bm{P}_{\bm{x}}$, which is orthogonal with respect to the Lorentzian inner product.
3.2 The GPHLVM Pullback Metric
Here, we consider the case where the mapping $f$ is defined as a GPHLVM. We derive the distribution of the Riemannian Jacobian, which we then use to compute the expected GPHLVM pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$. As for GPLVMs (Tosi et al., 2014), the joint distribution of the Jacobian $\tilde{\bm{J}}$ and observations $\bm{Y}=[\bm{y}_{1}\ldots\bm{y}_{N}]^{\mathsf{T}}\in\mathbb{R}^{N\times D_{y}}$ is given as,
@@ 3.2 The GPHLVM Pullback Metric
where $\bm{X}=[\bm{x}_{1}\ldots\bm{x}_{N}]^{\mathsf{T}}\in\mathbb{R}^{N\times(D_{x}+1)}$ with $\bm{x}_{n}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$, $\bm{K}_{X}\in\mathbb{R}^{N\times N}$ is the corresponding kernel matrix, $\bm{Y}_{d}$ denotes the $d$-th row of $\bm{Y}$, and we rely on the hyperbolic kernel derivatives,
$\partial k(\bm{X},\bm{x}^{*})$ $=\frac{\partial}{\partial\bm{z}}k(\bm{X},\bm{z})\Big{|}_{\bm{z}=\bm{x}^{*}},$ (16)
$\partial^{2}k(\bm{x}^{*},\bm{x}^{*})$ $=\frac{\partial^{2}}{\partial\bm{z}\,\partial\bm{x}}k(\bm{x},\bm{z})\Big{|}_{\bm{x}=\bm{z}=\bm{x}^{*}}\,.$ (17)
While in the Euclidean case the kernel derivatives are straightforward, the hyperbolic setting is more challenging, as will be discussed in Sec. 4. Conditioning on the observations $\bm{Y}_{d}\in\mathbb{R}^{N}$ results in the probability distribution of the Jacobian at $\bm{x}^{*}$ (Bishop, 2006, Chap 2.3),
@@
The expected pullback metric is computed using (14) as,
$\mathbb{E}[\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}]=\bm{P}_{\bm{x}^{*}}(\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J})\bm{P}_{\bm{x}^{*}}^{\mathsf{T}}\,,$ (22)
@@
3.3 Hyperbolic Pullback Geodesics
Having endowed the hyperbolic latent space with a pullback metric $\bm{G}^{\text{P},\mathcal{L}}$, we are now interested in computing shortest paths, i.e., geodesics, between latent hyperbolic points. While hyperbolic geodesics have a closed-form solution, they may not always align with the underlying data distribution (see Fig. 2). Instead, hyperbolic pullback geodesics naturally follow the data distribution, but require solving an optimization problem. Specifically, pullback geodesics are computed by minimizing the curve length, or equivalently the curve energy $E$ with respect to the pullback metric. When considering a discretized geodesic composed by a set of $M$ points $\bm{x}_{i}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$, this boils down to minimize,
$E=\sum_{i=0}^{M-2}\bm{v}_{i}^{\mathsf{T}}\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}_{i}}\bm{v}_{i}\,,\quad\text{with}\quad\bm{v}_{i}=\text{Log}_{\bm{x}_{i}}({\bm{x}_{i+1}}).$ (23)
While it is possible to iteratively optimize the curve points $\bm{x}_{i}$ directly, this often leads to uneven spacing among them. This issue can be addressed either by using a parametric curve on the manifold (Gousenbourger et al., 2014) and optimize its parameters instead of the points directly, or by incorporating the spline energy as a regularization factor in the optimization process (Heeren et al., 2018). We follow the latter. The spline energy is $E_{\text{spline}}\!\approx\!\sum_{i=1}^{M-2}\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x}_{i},\bar{\bm{x}}_{i})^{2}$, where $\bar{\bm{x}}_{i}=\text{Exp}_{\bm{x}_{i-1}}\left(\frac{1}{2}\,\text{Log}_{\bm{x}_{i-1}}(\bm{x}_{i+1})\right)$ is the geodesic midpoint between $\bm{x}_{i-1}$ and $\bm{x}_{i+1}$. The final optimization problem is,
@@ 3.3 Hyperbolic Pullback Geodesics
with $\lambda$ weighting the influence of the regularization. As the optimization parameters are Riemannian, we leverage Riemannian optimizers such as Riemannian Adam (Bécigneul & Ganea, 2019) to optimize (23). The computation of pullback geodesics is summarised by Algorithm 1 in App. C.
Figure 2: Left and middle: Euclidean and hyperbolic pullback metrics on a $\mathsf{C}$-shape trajectory ( ). The geodesics of the base manifold ( ) and the pullbacks geodesic ( ) are depicted on both Euclidean and hyperbolic spaces. Right: Curve energy along the geodesics on the Euclidean (top) and hyperbolic (bottom) cases. These include the base manifold geodesic ( ), the base manifold geodesic with energy (23) evaluated using the pullback metric ( ), and the pullback geodesic ( ).
@@ 4 Hyperbolic Kernel Derivatives
In this section, we discuss the challenges encountered when using autodiff tools, in particular PyTorch  (Paszke et al., 2019), to compute the derivatives of the hyperbolic SE kernels. Table 1 shows that autodiff-based kernel derivatives are approximately five times slower than custom analytic implementations in the 2D case. More importantly, in the 3D case, autodiff fails to compute the derivatives entirely.
Table 1: Average computation times in seconds over $100$ runs, each evaluating the pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ and its derivatives at a random point $\bm{x}$. PyTorch fails to compute the 3D kernel derivatives.
$k^{\mathbb{H}^{2}_{\mathcal{L}}}$ ($L=3000)$ (6) $k^{\mathbb{H}^{3}_{\mathcal{L}}}$ (5)
@@
4.1 2D Hyperbolic SE Kernel Derivatives
We compute the 2D hyperbolic SE kernel $k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{z})$ via the Monte Carlo approximation (6). The pullback metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ is computed using the kernel derivatives $\frac{\partial}{\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{x}^{*}}$ (16) and $\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$ (17). The only part of the kernel (6) that depends on the inputs $\bm{x},\bm{z}$ is the function $\Phi_{l}(\bm{x},\bm{z})$, whose derivatives are given by,
@@
The kernel derivatives (16)-(17) are thus straightforwardly determined from (25)-(26), and subsequently leveraged to compute $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ using (22). Minimizing the curve energy (23) to compute pullback geodesics additionally requires the derivative $\frac{\partial}{\partial\bm{x}^{*}}\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$. This, in turn, requires the derivative of the Jacobian of the covariance matrix $\frac{\partial}{\partial\bm{x}^{*}}\bm{\Sigma}_{J}$, which depends on the kernel derivatives $\frac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$, and $\frac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$. Again, the kernel derivatives are effectively determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$,
$\!\frac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}\Phi_{l}(\bm{x},\bm{z})\!$ $=\!\left[\!\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\phi_{l}(\bm{x}_{\mathcal{P}}\!)\!\right]\!\times_{2}\!\left[\!\dfrac{\partial}{\partial\bm{z}}\bar{\phi}_{l}(\bm{z}_{\mathcal{P}}\!)\!\right],$ (27)
$\!\frac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}\Phi_{l}(\bm{x},\bm{z})\!$ $=\!\left[\!\dfrac{\partial^{2}}{\partial\bm{z}^{2}}\bar{\phi}_{l}(\bm{z}_{\mathcal{P}}\!)\!\right]\!\times_{1}\!\left[\!\dfrac{\partial}{\partial\bm{x}}\phi_{l}(\bm{x}_{\mathcal{P}}\!)\!\right],$ (28)
@@
where $\bm{A}\times_{i}\bm{B}$ refers to n-mode product (Kolda & Bader, 2009b) which multiplies the $i$-th dimension of $\bm{A}$ with the second dimension of $\bm{B}$. If $\bm{B}$ is a vector, $\bm{A}$ is unsqueezed at dimension $i$ before the multiplication. The matrices $\frac{\partial^{2}}{\partial\bm{x}^{2}}\phi_{l}(\bm{x}_{\mathcal{P}}),\frac{\partial^{2}}{\partial\bm{z}^{2}}\phi_{l}(\bm{z}_{\mathcal{P}})\in\mathbb{R}^{3\times 3}$ in (27)-(28) are therefore interpreted as $\mathbb{R}^{3\times 1\times 3}$ and $\mathbb{R}^{1\times 3\times 3}$, respectively. The complete derivation of all derivatives is provided in App. B.
4.2 3D Hyperbolic SE Kernel Derivatives
@@
The 3D hyperbolic SE kernel is given by (5). In this section, we simplify its notation and write $k(\bm{x},\bm{z})=\frac{\rho}{s}e^{-\frac{\rho^{2}}{\nu}}$ by omitting the constant $\tau/C_{\infty}$, setting $\nu=2\kappa^{2}$, and introducing the helper variables $u=\langle\bm{x},\bm{z}\rangle_{\mathcal{L}}$ and $s=\sqrt{u^{2}-1}$. As in the 2D case, to obtain the expected pullback metric tensor (22), we compute the first two kernel derivatives,
$\dfrac{\partial}{\partial\bm{x}}\,k(\bm{x},\bm{z})$ $=g(u)\,\bm{G}^{\mathcal{L}}\bm{z}\,e^{-\frac{\rho^{2}}{\nu}}\,,$ (29)
$\dfrac{\partial^{2}}{\partial\bm{z}\partial{\bm{x}}}k(\bm{x},\bm{z})$ $=\left(h(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)\bm{G}^{\mathcal{L}}\right)e^{-\frac{\rho^{2}}{\nu}}\,,$ (30)